OpenStudy (anonymous):

What is the imaginary part of the complex number -11 – 3i?

OpenStudy (lgbasallote):

hint: the one with the i

hartnn (hartnn):

the imaginary part in x+iy is y does this help??

OpenStudy (anonymous):

a + i b a = real part y - imaginary

OpenStudy (anonymous):

here -11 – 3i? -11 = real -3 = imaginary